# CausalRAG: Integrating Causal Graphs into Retrieval-Augmented Generation

### v1

@@ 5.1 Experimental Setup
Parameters & Baselines. We set the default parameters for CausalRAG as $k=s=3$ and use the same $k$ value for GraphRAG’s community-based retrieval and regular RAG’s document retrieval to ensure fairness. We evaluate four RAG variants: regular RAG Lewis et al. (2021), GraphRAG with local and global search Edge et al. (2024), and our proposed CausalRAG. Local search retrieves from raw document graphs, making it ideal for passage-specific queries, while global search relies on graph community summaries for broader context understanding. We use GPT-4o-mini as the base LLM for all models.
Datasets. We evaluate on publicly available research papers from various domains, ranging from 100 to around 17,000 tokens, reflecting typical user-uploaded documents. An LLM will generate $n$ grounded questions per document, ensuring they are explicitly answerable. We use the Ragas framework Es et al. (2023) and assess models on three key metrics: answer faithfulness, context precision, and context recall. Faithfulness measures factual consistency on a scale from 0 to 1, with higher scores indicating better alignment with reference documents.
@@ 5.2 Performance Comparison
To comprehensively evaluate the effectiveness of CausalRAG, we conducted experiments on a diverse dataset comprising abstracts from 100 randomly selected research papers using OpenAlexPriem et al. (2022). All evaluations were performed using the OpenAlex public academic dataset, which spans multiple disciplines, including applied mathematics, art history, library science, and psychology, ensuring a comprehensive assessment across different domains. For each document, we constructed graph and generated $n=5$ questions per paper for assessing retrieval and response quality. As mentioned, we then compared CausalRAG’s performance against three baselines—regular RAG, GraphRAG-Local, and GraphRAG-Global—across three key metrics: answer faithfulness, context recall, and context precision.
Figure 3 presents the comparative results. Overall, CausalRAG consistently outperforms all other models across the three evaluation metrics, demonstrating its ability to generate more factually accurate responses while maintaining both high recall and precision.
Answer Faithfulness. This metric measures the extent to which the generated response aligns with the reference information, ensuring factual correctness. As shown in Figure 3, regular RAG achieves relatively strong performance in this metric, suggesting that despite its reliance on purely semantic retrieval, it still retrieves some relevant context. However, GraphRAG-Local still slightly outperforms regular RAG by leveraging structured graph-based retrieval locally, which supports its grounded answer. Naturally, GraphRAG-Global CausalRAG supported by high-level community summary performs worse in this metric. Lastly, CausalRAG holds its answer faithfulness at a good level by retrieving causally grounded context, reducing hallucinations, and ensuring responses are not only relevant but also justified by retrieved evidence.
Context Recall. Context recall evaluates how much of the correct reference information is retrieved during the RAG process. As expected, regular RAG exhibits high recall, given its broad retrieval approach, which tends to maximize the inclusion of potentially relevant content. GraphRAG-Local and GraphRAG-Global show a slight reduction in recall due to their community-based summarization, which, while improving retrieving process, sacrifices some context diversity. CausalRAG, by contrast, strikes a balance between precision and recall and its recall is still higher than regular RAG. It is ensured by that retrieved information is causally relevant, avoiding excessive retrieval of loosely related but non-causal information.
Context Precision. Precision is where CausalRAG demonstrates the most significant advantage. While regular RAG achieves relatively low precision—often retrieving semantically similar yet contextually irrelevant content—GraphRAG-Local and GraphRAG-Global significantly improve upon this by leveraging graph structures to organize knowledge more effectively. However, both still struggle with retrieving causally relevant content and are lower than CausalRAG on this metric. We can see CausalRAG still maintain a great precision, as it inherently filters retrieved information through causal graph.
Figure 3: Performance comparison of Regular RAG, GraphRAG-Global, GraphRAG-Local, and CausalRAG on the OpenAlex dataset across three key metrics: answer faithfulness, context recall, and context precision.
@@ 5.3 Case Study
Figure 4: Case Study comparing retrieving process between different RAG systems.
Figure 5: Performance comparison of four RAGs across documents of varying lengths (Abstract, Introduction, Full Paper) based on mean values of answer faithfulness, context recall, and context precision.
We also conducted a case study to evaluate the four RAG variants on a research paper with long text, but testing them on its abstract, introduction, and full text respectively. This approach ensures content consistency while varying document length (ranging from 255 to 16,475 tokens), allowing us to assess the scalability of CausalRAG.
For each of the three text materials, we constructed a separate graph, tested $n=20$ questions, and evaluated all four RAG variants across three key metrics. As mentioned earlier, we employed an LLM-based approach for graph construction, and the validity of these graphs—built from a business school marketing paper—was further verified by experts.
This case study also examines the retrieval differences among regular RAG, GraphRAG, and our proposed CausalRAG (as shown in Figure 4). In this example, a user uploads a long paper and asks: "How do different combinations of influence tactics impact the likelihood of winning a sales contract?" Expert validation confirms a clear answer—salespeople use influence tactics such as compliance-based and internalization-based strategies to attract buyers’ attention, which inherently increases their chances of securing a contract.
Analyzing the retrieval process, we observe that regular RAG retrieves semantically related content but fails to capture key information, leading to a vague and uninformative response. GraphRAG correctly retrieves relevant nodes like attention, but its community report process dilutes their importance by incorporating extraneous nodes, introducing bias despite factual grounding. In contrast, CausalRAG retrieves the initial relevant nodes and expands upon them by identifying causal paths essential for answering the query, ensuring a precise and causally grounded response.
While this example clarifies CausalRAG’s internal process, our full experimental results (Figure 5) demonstrate its scalability. CausalRAG consistently outperforms other RAG variants across different document lengths. Additionally, regular RAG performs well on shorter documents, even surpassing GraphRAG. However, as document length increases—such as with full papers—GraphRAG overtakes regular RAG, aligning with our analytical analysis.
CausalRAG exhibits robustness across varying context lengths by preserving access to discrete information while leveraging causal estimation for grounded retrieval. More importantly, an additional advantage—only partially reflected here—is that with a larger step size $s$, CausalRAG can trace long and complex causal relationships hidden in lengthy documents, generating well-founded responses—a capability that is challenging for existing RAG systems.
@@ 5.5 Conclusion and Future work
We introduced CausalRAG, a framework that integrates causality into RAG to enhance performance. Our analytical and experimental analysis identified key limitations of regular RAG—loss of contextual integrity, reliance on semantic similarity over causal relevance, and low precision. By leveraging causal graphs, CausalRAG retrieves causally grounded context, outperforming baseline RAGs across diverse text lengths and domains. Additionally, it improves LLM explainability, enhances answer groundedness, and reduces hallucinations in generative responses. Future work includes benchmarking CausalRAG on domain-specific tasks.
@@
While CausalRAG enhances retrieval effectiveness by integrating causality, it has certain limitations. First, constructing causal graphs from unstructured text relies on LLM-based extraction, which may introduce extra costs, particularly in complex or ambiguous cases. Second, the computational cost of expanding and analyzing causal paths increases with larger documents, potentially impacting retrieval efficiency in extreme cases with a large number of tokens.

### v2

@@ 5.1 Experimental Setup
Baselines. We evaluate five RAG variants: Regular RAG Lewis et al. (2021), GraphRAG with both local and global search Edge et al. (2024), HippoRAG2 Gutiérrez et al. (2025), and our proposed CausalRAG. Regular RAG serves as a standard baseline, relying solely on semantic similarity for retrieval. GraphRAG is a widely recognized framework that leverages graph community summaries, and we include both of its modes: the local version retrieves from raw document graphs and is well-suited for passage-level queries, while the global version summarizes graph communities to support broader context understanding. HippoRAG2 is a recent and competitive method that enhances retrieval by selecting seed nodes and ranking filtered triples for generation.
Datasets. While many open-domain QA benchmark datasets exist, most are designed for explicit fact retrieval (e.g., “When was Google founded?” “1998”) and target classic NLP tasks. These datasets often fall short in evaluating discourse-level understanding, such as querying the underlying ideas, logic, or narrative within a document—tasks that better reflect real-world needs. Although some reading comprehension datasets exist Rajpurkar et al. (2018), their answers are typically short entities and designed for NLP models rather than RAG systems.
To more effectively evaluate RAGs in knowledge-intensive tasks, recent research calls for datasets that require higher-level discourse understanding, such as those based on podcasts, news articles, or Wikipedia Edge et al. (2024); Gutiérrez et al. (2025). Following this direction, we use academic papers sampled from the OpenAlex dataset Priem et al. (2022a). Dataset statistics are provided in Table 1. For each document, we use an LLM to generate $n=5$ grounded questions, ensuring they are explicitly answerable (see Appendix B for examples).
Metrics & implementation details. We use the Ragas evaluation framework Es et al. (2023) to assess all models on three metrics: answer faithfulness, context precision, and context recall. The definitions of context precision and context recall are provided in the previous section. Answer faithfulness measures factual consistency on a scale from 0 to 100, with higher scores indicating closer alignment with reference documents. We use GPT-4o-mini as the base LLM for all frameworks. We set the parameters $k=s=3$ for CausalRAG, and use the same $k$ value for GraphRAG’s community-based retrieval, Regular RAG’s document retrieval, and HippoRAG2’s triple-based retrieval to ensure a fair comparison.
Figure 4: Case Study – A user uploads a long paper and asks a related question. This figure compares Regular RAG, GraphRAG, and CausalRAG by analyzing their retrieval processes. It highlights the drawbacks of semantic and graph-based retrieval and shows how causal reasoning in CausalRAG leads to more robust and precise results.
Figure 5: Case Study – A follow-up experiment evaluates the RAGs in the previous case using three versions of the same paper. Graph-based methods improve with length, while CausalRAG remains consistently robust.
@@ 5.2 Performance Comparison
Figure 3 presents the main experimental results comparing five RAG frameworks across three evaluation metrics: answer faithfulness, context recall, and context precision. The results reveal distinct patterns in how different retrieval strategies affect generation quality. Below, we summarize the key findings and underlying trade-offs.
Causality delivers the most balanced and accurate retrieval. The model incorporating causal reasoning achieves the highest performance across all three metrics, with a context–precision score of $92.86$ and a faithfulness score of $78.00$, while maintaining competitive recall ($49.46$). By grounding retrieval in explicit causal paths, it filters out semantically similar but logically irrelevant content, thereby reducing superficial answers and improving the alignment and coherence between the retrieved context and the user query.
Graph-level structuring improves precision but introduces a recall–coverage trade-off. GraphRAG’s two operating modes illustrate this tension. Local search, which queries the raw knowledge graph, yields strong precision ($89.18$) but discards many supporting passages, dropping recall to $41.54$. Global search restores some of that missing evidence by summarizing community subgraphs, yet the broader scope simultaneously re-introduces off-topic material, pulling precision down to $66.67$ and faithfulness below $55.3$. These results corroborate Edge et al. (2024)’s observation that finer-grained community answers excel at factual grounding, whereas higher-level summaries aid breadth but dilute specificity.
Entity-centric multi-stage KGs narrow the gap but still overlook discourse context. HippoRAG2 augments Personalized PageRank with LLM-filtered triples and passage nodes, boosting precision to $73.72$ and faithfulness to $67.36$ without sacrificing recall relative to Regular RAG. Nevertheless, its reliance on entity extraction leaves many context-rich sentences unlinked to the query, a limitation they noted as concept–context trade-off and need for deeper contextualizationGutiérrez et al. (2025). By contrast, CausalRAG’s path-based expansion integrates both entities and their causal explanations, yielding a retrieval set that is simultaneously broader in coverage and more tightly aligned with the question, thereby establishing a new possibility for knowledge-intensive RAG tasks.
@@ 5.3 Case Study
Qualitative exemplar. Figure 4 traces the end-to-end behaviour of the three systems on a single question drawn from a marketing paper: “How do different combinations of influence tactics impact the likelihood of winning a sales contract?” The semantic baseline retrieves passages that mention “influence tactics” but stop short of establishing how those tactics translate into contract wins; its answer therefore echoes the query without providing the missing mechanism. Graph-based retrieval narrows the search space by clustering entities such as buyer attention and sales contract award, yet the community summary also injects peripheral phrases, leading the model to an over-generalised explanation that only partially matches the ground truth. By contrast, the causality-driven pipeline first aligns the query with a causal pathway—Influence Tactics $\rightarrow$ Buyer Attention $\rightarrow$ Contract Award. Because that pathway is preserved through retrieval and summarisation, the generated answer pinpoints the mediating role of buyer attention and specifies which tactic pairs are most effective, fully satisfying expert judgment.
Length-controlled follow-up. To test whether these qualitative differences persist when context grows, we replicate the experiment on three versions of the same paper—abstract (250 tokens), introduction (1k), and full text (16k)—and score each system on the composite of faithfulness, precision, and recall (Figure 5). Graph-level methods benefit from additional material: their aggregate score climbs from 49.98/62.29 on the abstract to 76.37/84.27 on the full paper, confirming that community-based indexing scales gracefully with document length. Yet the causality-oriented approach remains consistently ahead, posting 72.43 on the abstract, 74.86 on the introduction, and 91.69 on the full text. This steadiness indicates that causal expansion recovers the right evidence even when local term overlap is sparse (short documents) and still filters noise when semantic matches proliferate (long documents).
Emerging pattern. Taken together, the case study and its quantitative extension suggest a general rule: graph structure raises precision by adding relational cues, but explicit causal alignment is required to preserve both precision and recall across scales. Semantic search alone misses latent mechanisms; graph aggregation alone retains residual noise. Embedding causal constraints into retrieval therefore offers a principled path to robust performance on knowledge-intensive tasks, regardless of document length or query granularity.
@@ 5.5 Conclusion and Future Work
We introduced CausalRAG, a novel framework that integrates causal reasoning into retrieval-augmented generation to address core limitations of existing RAG systems. By leveraging causality, CausalRAG retrieves context that is not only relevant but also causally grounded—enhancing generation quality, reducing hallucinations, and improving answer fidelity.
Future work may extend CausalRAG in several directions. A key area is testing scalability in long-context scenarios, which requires constructing and processing numerous high-dimensional graphs, each derived from large-scale documents spanning millions of tokens. While current practices often merge documents into a single graph or use smaller graphs due to data quality and computational constraints, this direction is critical for understanding RAG scalability. In addition to improving performance, evaluating efficiency and generation latency remains important for practical deployment.
@@
While CausalRAG improves retrieval effectiveness through causal reasoning, it has certain limitations. First, the approach relies on the internal knowledge of LLMs to construct graphs and identify causal relationships. Although both capabilities have been actively studied, emerging domain-specific knowledge—such as in medicine or law—may still constrain its effectiveness in specialized contexts. Second, identifying causal paths during inference also requires additional LLM calls, introducing extra computational costs that may limit the performance in real-world deployments.
Appendix
Figure 7: LLM prompts for causal discovery and causal summary
Appendix A LLM Prompts Used
The full prompts used in the CausalRAG framework for causal discovery and causal summarization (Figure 7), as illustrated in Figure 2. The prompts are adapted from Edge et al. (2024).
Appendix B OpenAlex Dataset Example
This is an exmaple (see Table 2) in the dataset Priem et al. (2022b) along with generated questions.
Dataset OpenAlex Priem et al. (2022b)
Discipline Atomic Physics
Author Thom H. Dunning (1989)
Text In the past, basis sets for correlated molecular calculations were largely taken from single-configuration calculations. Recently, Almlöf, Taylor, and co-workers showed that atomic natural orbitals (ANOs) provide an excellent description of correlation effects. Here we report a careful oxygen-atom study establishing that compact primitive Gaussian functions effectively describe correlation when their exponents are optimized. Tests on oxygen-containing molecules indicate these functions perform as well as the ANO sets of Almlöf and Taylor. Guided by the oxygen results, basis sets were developed for all first-row atoms (B–Ne) and hydrogen. Incremental energy lowerings due to correlating functions fall into distinct groups, leading to the concept of consistent sets. The most accurate set, [5s 4p 3d 2f 1g], consistently yields 99% of the correlation energy obtained with the next larger set, even though the latter contains 50% more primitives and twice as many polarization functions. For boron, this set recovers 94–97% of the total (HF+1+2) correlation energy.
Questions 1. What recent findings support the use of atomic natural orbitals in molecular calculations? 2. How do compact primitive Gaussian functions contribute to describing correlation effects in oxygen? 3. Why is exponent optimization important in Gaussian-based calculations? 4. How do the new first-row basis sets compare in energy lowering from correlation effects? 5. What accuracy (%) is achieved for boron with the most compact set, and how does this relate to the number of polarization functions?
Table 2: Example document metadata, full-text excerpt, and evaluation questions used in our study.